# Enhancing Sentiment Analysis in Bengali Texts: A Hybrid Approach Using Lexicon-Based Algorithm and Pretrained Language Model Bangla-BERT

### v1

@@ III Methodology
    The Bangla Sentiment Polarity Score (BSPS) is an advanced sentiment analysis algorithm is specifically designed for processing Bengali text which have semantic meaning and dialect complexities. Leveraging a combination of LDD and linguistic rules, BSPS evaluates the sentiment of sentences analyzing key components such as positive and negative words, negation terms, extreme modifiers, phrase initiator etc. and by generating score for each review. The system is adept at handling the nuances of the Bengali language, including word order and the impact of conjunctions, phrases, and negations on sentiment. By utilizing a comprehensive lexicon, BSPS attempts to provide an accurate sentiment score, helping to understand the underlying emotional tone of Bengali texts.
@@
        Initialization: The algorithm initializes a sentiment score (score = 0) and flags for tracking various conditions (e.g., neg-flag, pos-word-flag, etc.).
@@
        Processing Each Word: The algorithm iterates through each word in the filtered sentence and processes it based on whether it’s found in any of the predefined lexicons and lists (positive words, negative words, Direct Negation Words, Extreme Words etc.).
@@
        Handling ”AND” Words: If an ”and-word” (like [Uncaptioned image] [and]) is encountered, it triggers the and-flag. The algorithm looks at the previous and next words and applies a special rule for combining their sentiments (e.g., two positive words joined by ”and” might amplify the sentiment).
@@
        Handling Phrases: If a phrase-initial word (like [Uncaptioned image] [that much]) is found, it triggers the phrase-flag. The algorithm then checks whether the word after the phrase initial is positive or negative and applies context-specific rules (e.g. [Uncaptioned image] [that much good] might indicate an intensified positive sentiment).
@@
        Handling Extreme Words: If an extreme word is encountered, it triggers the extreme-flag. Extreme words amplify the sentiment score by a factor (e.g., [Uncaptioned image] [very good] becomes stronger than just ”[Uncaptioned image]” [good]). The algorithm checks if the extreme word is associated with a positive or negative word, applying the appropriate multiplier (e.g., [Uncaptioned image] [very good] → score * 1.6).
@@
        Handling Negation: When a negation word (like [Uncaptioned image] [no or not]) is found, it triggers the neg-flag. The algorithm applies rules to reverse the sentiment of the associated word.
@@
        Final Calculation: While processing all the words, the algorithm computes and updates the sentiment score based on each individual word it encounters, adjusted by rules for negation, conjunctions, and extremity.
@@
        Example-01: Given the sentence [Uncaptioned image] [The food was not very good and delicious], here’s how the algorithm processes it:
@@
        Example-02: Given the sentence [Uncaptioned image] [So good that it can’t be believed], here’s how the algorithm processes it:
@@
        When the algorithm encounters the word ”[Uncaptioned image]”, it activates the phrase-flag. As a result, the direct negation word ”[Uncaptioned image]” does not reverse the sentiment but instead significantly amplifies the sentiment score. It makes sense as the sentence [Uncaptioned image] [So good that it can’t be believed] surely implies that the product is very very good.
@@
        In this case, we can see that, the sentence [Uncaptioned image] [The bag is not very bad] essentially conveys the meaning of [Uncaptioned image] [The bag is somewhat good]. Traditional rule-based algorithms typically fail to account for this contextual interpretation. They would simply calculate [Uncaptioned image] [very bad], and then negate it due to the word ”[Uncaptioned image]” [not], resulting in [Uncaptioned image] [the bag is very good], which misrepresents the actual meaning of the sentence. However, our Bangla Sentiment Polarity Score(BSPS) algorithm carefully considers the subtle difference in meaning and incorporates this contextual nuance which is a coherent feature of Bengali language to produce the correct sentiment.
    Figure 2: BSPS Algorithm
    Figure 3: BSPS Algorithm cont.
    Figure 4: BSPS Algorithm cont.
    Figure 5: BSPS Algorithm cont.
@@
    Once we had generated sentiment scores for each review in our dataset using the Bengali Sentiment Polarity Scoring (BSPS) algorithm, we undertook a thorough analysis of these scores. The raw sentiment scores were carefully evaluated to determine the emotional tone conveyed by each review. To ensure consistency and comparability across the reviews, we applied a normalization process to the scores.
@@ IV Results
To present the results of our work, we start by analyzing and evaluating the intermediate outcomes of the classification processes in both the BSPS and BanglaBERT models. In this stage, we predicted sentiments (limited to positive and negative as per the original dataset) for the entire dataset using: 1) the BSPS algorithm, which provides both sentiment labels and sentiment polarity scores, and 2) the pre-trained BanglaBERT model, which outputs probability values from which we inferred the sentiments for the entire dataset. We compare the performance of these two methods in predicting the initial (positive and negative) sentiments. Following this, we examine the evaluation results and compare the two approaches used in our study: 1) sentiment classification with BSPS and evaluation with BanglaBERT, and 2) both sentiment classification and evaluation using BanglaBERT.
@@
    Performance of the BSPS Algorithm: In our first approach of the BSPS algorithm for sentiment classification, the model achieved an impressive accuracy of 93%, demonstrating its effectiveness in distinguishing between positive and negative sentiments. The algorithm produced both sentiment labels and polarity scores for the entire dataset, ensuring alignment with the original dataset’s classification. The BSPS algorithm’s strong performance metrics underscore its reliability and potential as an advanced tool for sentiment analysis.
@@
    Performance of the BanglaBERT (intermediate): In the intermediate phase of our study, we applied the BanglaBERT model for sentiment classification, focusing on positive and negative sentiment labels. The model achieved an accuracy of 88%, showcasing its ability to correctly classify sentiment in a substantial portion of the dataset. BanglaBERT produced precise sentiment labels and corresponding polarity scores, which were analyzed to evaluate its effectiveness. Despite a strong performance, the metrics indicate potential areas for improvement compared to the BSPS algorithm. Overall, BanglaBERT demonstrated solid performance, making it a valuable tool for sentiment analysis in the Bangla language.
@@
    Although BanglaBERT is a powerful tool, it may not be as good at handling the subtle distinctions between closely related sentiment categories as BSPS. The following figure[Fig.11] illustrates the comparison of the two models’ performance in the intermediate level:
@@
    The results indicate that BanglaBERT (classification) + BanglaBERT (evaluation) underperforms compared to BSPS (classification) + BanglaBERT (evaluation) across all evaluation metrics, including accuracy (89% vs. 79%), weighted precision (0.89 vs. 0.69), weighted recall (0.89 vs. 0.79), and weighted F1-score (0.89 vs. 0.73). This suggests that the combination of using the BSPS algorithm for classification followed by BanglaBERT for evaluation yields better performance overall. For BSPS (classification) + BanglaBERT (evaluation): The best performance was achieved with learning rate = 1.4e-5 and batch size = 32. For BanglaBERT (classification) + BanglaBERT (evaluation): The best performance was achieved with learning rate = 1.8e-4 and batch size = 16. The BSPS algorithm’s lower learning rate (1.4e-5) and larger batch size (32) likely contribute to more stable training and more accurate sentiment predictions. In contrast, the BanglaBERT-only pipeline, with a higher learning rate (1.8e-4) and smaller batch size (16), may have struggled with effective learning, which impacted its predictive performance. This comparison underscores the advantage of leveraging the BSPS algorithm for the initial classification step while using BanglaBERT for evaluation, although further optimization of BanglaBERT’s hyperparameters could improve its results when used alone.
@@ V Conclusion
This highlights the effectiveness of the hybrid approach, where the BSPS algorithm efficiently pre-processes and categorizes reviews into broader sentiment categories, setting a strong foundation for the subsequent detailed analysis with BanglaBERT. The integration leads to better performance, quicker learning, and more accurate results. Pipeline 2, while effective for binary classification tasks, showed limitations in handling complex multi-category sentiment analysis and may need further optimization for enhanced performance in nuanced sentiment classifications.
@@ VI Future Work
Future improvements could involve experimenting with more sophisticated pre-trained models like mBERT or XLM-R, which might offer better performance by leveraging multilingual understanding and broader linguistic nuances. Additionally, integrating the outputs of Pipeline 1 and Pipeline 2 through ensemble techniques could create a hybrid model that harnesses the strengths of both approaches, leading to more robust sentiment analysis across different types of data and scenarios.

### v2

@@ III Methodology
    The Bangla Sentiment Polarity Score (BSPS) algorithm (Algorithm 1) is an advanced sentiment analysis algorithm is specifically designed for processing Bengali text which have semantic meaning and dialect complexities. Leveraging a combination of LDD and linguistic rules, Algorithm 1 evaluates the sentiment of sentences analyzing key components such as positive and negative words, negation terms, extreme modifiers, phrase initiator etc. and by generating score for each review. The system is adept at handling the nuances of the Bengali language, including word order and the impact of conjunctions, phrases, and negations on sentiment. By utilizing a comprehensive lexicon, Algorithm 1 attempts to provide an accurate sentiment score, helping to understand the underlying emotional tone of Bengali texts.
@@
        Initialization: Algorithm 1 initializes a sentiment score (score = 0) and flags for tracking various conditions (e.g., neg-flag, pos-word-flag, etc.).
@@
        Processing Each Word: Algorithm 1 iterates through each word in the filtered sentence and processes it based on whether it’s found in any of the predefined lexicons and lists (positive words, negative words, Direct Negation Words, Extreme Words etc.).
@@
        Handling ”AND” Words: If an ”and-word” (like [Uncaptioned image] [and]) is encountered, it triggers the and-flag. Algorithm 1 looks at the previous and next words and applies a special rule for combining their sentiments (e.g., two positive words joined by ”and” might amplify the sentiment).
@@
        Handling Phrases: If a phrase-initial word (like [Uncaptioned image] [that much]) is found, it triggers the phrase-flag. Algorithm 1 then checks whether the word after the phrase initial is positive or negative and applies context-specific rules (e.g. [Uncaptioned image] [that much good] might indicate an intensified positive sentiment).
@@
        Handling Extreme Words: If an extreme word is encountered, it triggers the extreme-flag. Extreme words amplify the sentiment score by a factor (e.g., [Uncaptioned image] [very good] becomes stronger than just ”[Uncaptioned image]” [good]). Algorithm 1 checks if the extreme word is associated with a positive or negative word, applying the appropriate multiplier (e.g., [Uncaptioned image] [very good] → score * 1.6).
@@
        Handling Negation: When a negation word (like [Uncaptioned image] [no or not]) is found, it triggers the neg-flag. Algorithm 1 applies rules to reverse the sentiment of the associated word.
@@
        Final Calculation: While processing all the words, Algorithm 1 computes and updates the sentiment score based on each individual word it encounters, adjusted by rules for negation, conjunctions, and extremity.
@@
        Example-01: Given the sentence [Uncaptioned image] [The food was not very good and delicious], here is how Algorithm 1 processes it:
@@
        Example-02: Given the sentence [Uncaptioned image] [So good that it can’t be believed], here’s how Algorithm 1 processes it:
@@
        When Algorithm 1 encounters the word ”[Uncaptioned image]”, it activates the phrase-flag. As a result, the direct negation word ”[Uncaptioned image]” does not reverse the sentiment but instead significantly amplifies the sentiment score. It makes sense as the sentence [Uncaptioned image] [So good that it can’t be believed] surely implies that the product is very very good.
@@
        In this case, we can see that, the sentence [Uncaptioned image] [The bag is not very bad] essentially conveys the meaning of [Uncaptioned image] [The bag is somewhat good]. Traditional rule-based algorithms typically did not account for this contextual interpretation. They would simply calculate [Uncaptioned image] [very bad], and then negate it due to the word ”[Uncaptioned image]” [not], resulting in [Uncaptioned image] [the bag is very good], which misrepresents the actual meaning of the sentence. However, our Algorithm 1 carefully considers the subtle difference in meaning and incorporates this contextual nuance which is a coherent feature of Bengali language to produce the correct sentiment.
    Algorithm 1 BSPS ALGORITHM
    1:Input: Sentence as a string
    2:Output: Sentiment score (float)
    3:Tokenize the sentence:
[…]
    6:Initialize: $score\leftarrow 0$
[…]
    8:for each word in filtered_tokens do
    9:     if word in and_word then
    10:         $and\_flag\leftarrow\text{True}$
    11:     else if word in phrase_initial then
    12:         $phrase\_flag\leftarrow\text{True}$
    13:     else if word in extreme_words then
    14:         $extreme\_flag\leftarrow\text{True}$
    15:     else if word in direct_negation then
    16:         if any flag is True then
    17:              Adjust score based on flags
    18:              Reset flags
    19:         else
    20:              $neg\_flag\leftarrow\text{True}$
    21:         end if
    22:     else if word in positive_lexicon then
    23:         if any flag is True then
    24:              Adjust score based on flags
    25:              Reset flags
    26:         else
    27:              $pos\_word\_flag\leftarrow\text{True}$
    28:              Update score
    29:         end if
    30:     else if word in negative_lexicon then
    31:         if any flag is True then
    32:              Adjust score based on flags
    33:              Reset flags
    34:         else
    35:              $neg\_word\_flag\leftarrow\text{True}$
    36:              Update score
    37:         end if
    38:     end if
    39:end for
    40:return score
@@
    Once we had generated sentiment scores for each review in our dataset using Algorithm 1, we undertook a thorough analysis of these scores. The raw sentiment scores were carefully evaluated to determine the emotional tone conveyed by each review. To ensure consistency and comparability across the reviews, we applied a normalization process to the scores.
@@
        Hyper-parameters: During fine-tuning, we tested 10 different sets of learning rates (ranging from 1e-5 to 1e-3) and batch sizes (8, 16, and 32) to identify the optimal model performance.
@@
        Training Process: The model was fine-tuned over three epochs. We used the ”Optuna” optimizer along with a dynamic learning rate scheduler and different batch sizes. The goal was for the model to effectively map input reviews to one of the nine sentiment categories.
@@ IV Results
To present the results of our work, we start by analyzing and evaluating the intermediate outcomes of the classification processes in both the BSPS and BanglaBERT models. In this stage, we predicted sentiments (limited to positive and negative as per the original dataset) for the entire dataset using: 1) the BSPS algorithm, which provides both sentiment labels and sentiment polarity scores, and 2) the pre-trained BanglaBERT model, which outputs probability values from which we inferred the sentiments for the entire dataset. We compare the performance of these two methods in predicting initial (positive and negative) sentiments. Following this, we examine the evaluation results and compare the two approaches used in our study: 1) sentiment classification with BSPS and evaluation with BanglaBERT, and 2) both sentiment classification and evaluation using BanglaBERT.
@@
    Performance of the BSPS Algorithm: In our first approach of the BSPS algorithm for sentiment classification, the model achieved an impressive accuracy of 93%, demonstrating its effectiveness in distinguishing between positive and negative sentiments. The algorithm produced both sentiment labels and polarity scores for the entire dataset, ensuring alignment with the original dataset’s classification. The BSPS algorithm performs well, showing that it’s both reliable and has the potential to be a valuable tool for analyzing sentiment in text.
@@
    Performance of the BanglaBERT (intermediate): In the intermediate phase of our study, we applied the BanglaBERT model for sentiment classification, focusing on positive and negative sentiment labels. The model achieved an accuracy of 88%, showcasing its ability to correctly classify sentiment in a substantial portion of the dataset. Despite having some difficulties with unbalanced data, BanglaBERT produced precise sentiment labels and the corresponding polarity scores, which were analyzed to evaluate its effectiveness. Despite strong performance, the metrics indicate potential areas for improvement compared to the BSPS algorithm. Overall, BanglaBERT demonstrated solid performance, making it a valuable tool for sentiment analysis in the Bangla language.
@@
    Although BanglaBERT is a powerful tool, it may not be as good at handling the subtle distinctions between closely related sentiment categories as BSPS. The following figure[Fig.11] Illustrates the performance comparison between the two models at the intermediate level:
@@
    The results indicate that BanglaBERT (classification) combined with BanglaBERT (evaluation) underperforms compared to BSPS (classification) + BanglaBERT (evaluation) across all evaluation metrics, including the accuracy (89% vs. 79%), weighted precision (0.89 vs. 0.69), weighted recall (0.89 vs. 0.79), and weighted F1-score (0.89 vs. 0.73). This suggests that the combination of using the BSPS algorithm for classification followed by BanglaBERT for evaluation yields better overall performance. For BSPS (classification) + BanglaBERT (evaluation): The best performance was achieved with the learning rate = 1.4e-5 and batch size = 32. For BanglaBERT (classification) + BanglaBERT (evaluation): The best performance was achieved with the learning rate = 1.8e-4 and batch size = 16. The lower learning rate (1.4e-5) of the BSPS algorithm and the larger batch size (32) likely contribute to more stable training and more accurate sentiment predictions. In contrast, the BanglaBERT-only pipeline, with a higher learning rate (1.8e-4) and smaller batch size (16), may have struggled with effective learning, which impacted its predictive performance. This comparison underscores the advantage of leveraging the BSPS algorithm for the initial classification step while using BanglaBERT for evaluation, although further optimization of BanglaBERT’s hyperparameters could improve its results when used alone.
@@ V Conclusion
This highlights the effectiveness of the hybrid approach, where the BSPS algorithm efficiently preprocesses and categorizes reviews into broader sentiment categories, setting a strong foundation for the subsequent detailed analysis with BanglaBERT. Integration leads to better performance, faster learning, and more accurate results. Pipeline 2, while effective for binary classification tasks, showed limitations in handling complex multicategory sentiment analysis and may need further optimization for enhanced performance in nuanced sentiment classifications.
@@ VI Future Work
Future improvements could involve experimenting with more sophisticated pre-trained models like mBERT or XLM-R, which might offer better performance by leveraging multilingual understanding and broader linguistic nuances. Furthermore, integrating the output of Pipeline 1 and Pipeline 2 through ensemble techniques could create a hybrid model that harnesses the strengths of both approaches, leading to more robust sentiment analysis across different types of data and scenarios.